OpenStudy (anonymous):

1 over 4 (2x − 14) = 4

OpenStudy (anonymous):

first you have this: 1/4(2x-14)=4 then you do this: 4*(1/4)(2x-14) = 4*4 then this: 2x-14 = 16 now this: 2x-14 + 14 = 16 +14 after that you get this: 2x = 30 and then you have this: x = 15

OpenStudy (anonymous):

thanks!